# Semantic Structure-Aware Generative Attacks for Enhanced Adversarial Transferability

### v3

@@ C.2 Implementation details
We train the perturbation generator for one epoch on the ImageNet-1K [55] dataset using the Adam [32] optimizer $\beta=(0.5,0.99)$. We set the learning rate $lr=2e^{-4}$.
@@
Dataset ImageNet-1K [55] CUB-200-2011 [67] Stanford Cars [33] FGVC Aircraft [45]
@@
Resolution 224$\times$224 448$\times$448 448$\times$448 448$\times$448
We describe the statistics of the datasets used for training and evaluation in Table C.2. Note that we do not use the training sets from CUB-200-2011 [67], Stanford Cars [33], or FGVC Aircraft [45] for the strict black-box cross-domain setting.
@@
Since our approach only involves computational overheads during the training of the perturbation generator, there is no inference time overhead. During training, we describe in the table on the right that the time for a single forward pass with a batch size of 1 incurs an additional +12.18 (ms) relative to the baseline [86] time of 44.72 (ms), and an additional +28.31 (MB) in memory relative to the baseline memory requirement of 1,404.13 (MB), which are averaged over 1,000 iterations (See Table S4). We note that the training was performed on a single NVIDIA RTX A6000 GPU.
@@ C.3 Additional quantitative results
While we performed experiments against the VGG-16 surrogate model for fair comparison with previous works, we also provide, in Tables S5 and S6, our improvements when trained against other surrogate models such as VGG-19, ResNet-152, and DenseNet-169 as practiced in [86, 82, 83]. Ours added to the baseline trained against all three surrogate models across models and domains, except for cross-model against Dense169, consistently enhances the attack transferability. As with our results against VGG-16 in Tables 1 and 2, our method effectively boosts the attack capacity regardless of the type of surrogate model used for training the generator. We believe the slight increase in cross-model average when using DenseNet-169 as the surrogate model is driven mainly by pronounced gains on a few architectures (e.g. MNASNet and DeiT) while most other models also benefited, albeit to a lesser extent. In the cross-domain evaluation, adversarial examples generated with DenseNet-169 consistently deliver a substantial average performance boost, underscoring its effectiveness across differing data distributions.
In qualitative comparisons across different surrogate models in Fig. S4, our Grad-CAM [56] visualizations reveal activation patterns that depart significantly from the baseline. Rather than highlighting only the main object region, our method further amplifies those top responses and draws out additional high-sensitivity areas that the baseline misses. When we step through the stages of adversarial noise generation, we see that our approach consistently places perturbations along object edges and contours. By focusing noise on these shared, model-agnostic features instead of scattering it elsewhere, our method not only seeks to align the generated adversarial noise with the most semantically meaningful regions but also achieves stronger transferability.
Table S5: Additional quantitative cross-model transferability results. We report the average improvement margins of our method added to the baseline [86] using different surrogate models (VGG-19, Res152, Dense169). We report the improvements ($\Delta$p) with ours relative to the baseline. Better averaged results are marked in boldface.
@@
BIA [86] w/ Ours Acc. ($\Delta$p) $\downarrow$ -1.49 +0.88 -1.70 -3.49 -4.04 -1.07 +1.06 -2.31 +0.46 -1.23 -4.40 +0.39 -0.24 -0.16 -6.38 -0.17 -0.32 +0.10 -0.55 -2.22 -2.29 -1.72 -1.40
ASR ($\Delta$p) $\uparrow$ +1.93 -1.12 +2.17 +4.43 +5.18 +1.34 -1.43 +3.98 -0.65 +1.77 +5.18 -0.44 +0.35 +0.20 +7.63 +0.25 +0.32 -0.24 +0.75 +2.58 +2.75 +1.91 +1.77
FR ($\Delta$p) $\uparrow$ +1.72 -0.89 +1.97 +3.95 +4.68 +1.24 -1.19 +2.89 -0.55 +1.42 +4.77 -0.37 +0.42 +0.15 +7.08 +0.09 +0.33 -0.33 +0.72 +2.48 +2.67 +1.95 +1.60
ACR ($\Delta$p) $\downarrow$ -0.23 +0.11 -0.35 -0.56 -0.41 -0.13 +0.34 -0.18 -0.20 -0.29 -0.82 0.24 +0.14 -0.02 -1.45 +0.18 -0.33 -0.24 +0.06 -0.93 -0.68 -0.75 -0.30
@@
BIA [86] w/ Ours Acc. ($\Delta$p) $\downarrow$ -0.61 -0.99 -2.96 -0.38 -5.13 -7.36 -5.35 -0.71 -3.87 -0.89 -6.17 -3.75 -0.12 -0.84 +0.67 -0.94 +0.20 +0.33 -2.37 -4.19 -9.23 -0.46 -2.51
ASR ($\Delta$p) $\uparrow$ +0.78 +1.28 +3.76 +0.39 +6.36 +9.09 +7.37 +1.32 +4.75 +1.24 7.16 +4.70 +0.18 +1.13 -0.97 +0.92 -0.24 -0.32 +3.03 +5.09 +11.10 +0.48 +3.12
FR ($\Delta$p) $\uparrow$ +0.78 +1.09 +3.23 +0.27 +5.70 +8.14 +6.29 +0.97 +4.37 +0.86 +6.79 +4.18 +0.17 +1.35 -1.04 +1.08 -0.28 -0.55 +2.96 +5.01 +10.83 +0.38 +2.84
ACR ($\Delta$p) $\downarrow$ -0.10 +0.01 -0.69 -0.35 -1.21 -1.26 -1.33 +0.06 -0.50 -0.29 -1.56 -0.68 +0.07 +0.19 -0.50 -1.06 +0.02 +0.35 -0.37 -0.96 -2.64 -0.34 -0.60
@@
BIA [86] w/ Ours Acc. ($\Delta$p) $\downarrow$ -2.25 -7.65 -2.24 -2.76 -12.99 +4.65 +10.36 +2.34 +31.22 -6.50 +1.19 -6.71 -0.62 -0.73 +11.05 +8.35 +0.38 +0.58 +3.63 -2.67 +11.51 -7.61 +1.48
ASR ($\Delta$p) $\uparrow$ +2.77 +9.41 +2.77 +3.38 +16.23 -5.34 -16.99 -6.14 -37.55 +9.68 -1.55 +8.22 +0.64 +0.70 -13.29 -9.27 -0.70 -0.77 -4.34 +3.20 -14.21 +8.67 -2.02
FR ($\Delta$p) $\uparrow$ +2.41 +8.48 +2.32 +2.88 +13.97 -4.84 -12.84 -4.23 -33.64 +7.93 -1.23 +7.24 +0.76 +0.75 -12.09 -9.48 -0.70 -0.65 -4.29 +2.99 -12.99 +8.44 -1.76
ACR ($\Delta$p) $\downarrow$ -0.73 -1.65 -0.71 -0.83 -2.64 +1.17 +1.78 +0.18 +6.26 -0.93 -0.40 -1.75 -0.54 -0.80 +2.23 +4.10 -1.31 +0.16 +1.46 -0.80 +2.00 -2.56 +0.17
Table S6: Additional quantitative cross-domain transferability results. We report the average improvement margins of our method added to each baseline, averaged over three models for each domain using different surrogate models (VGG-19, Res152, Dense169). We report the improvements ($\Delta$p) with ours relative to the baseline [86]. Better averaged results are marked in boldface.
Cross-domain CUB-200-2011 [67] Stanford Cars [33] FGVC Aircraft [45] Avg. Acc. (%)
@@
Figure S4: Comparison of the crafted AEs, Grad-CAMs, intermediate feature activation maps from the baseline, and with ours (columns 1–4, respectively). We present the qualitative results from training against other surrogate models: VGG-19 (Left), ResNet-152 (Center), and DenseNet-169 (Right) as commonly compared in existing generative attacks [86, 1, 82, 83]. The correct label and attacked prediction results are marked in green and white, respectively.
@@
DeepLabV3+ [11] SegFormer [78] Faster R-CNN [20] DETR [8]
@@
To further verify that our proposed method is compatible with the baseline [86] shown to perform best at the selected mid-level layer of the VGG-16 surrogate model (i.e., Maxpooling.3), we conducted an ablation study on the mid-level layer in Fig. S5. Across domain, model, and two tasks (SS and OD), we observe that our method added to the baseline still maintains the best strength of the attack at the selected mid-level layer, Maxpooling.3, compared to the other early or late layers.
@@ C.4 Additional qualitative results
Figure S7: Additional qualitative results. Our semantic structure-aware attack successfully guides the generator to focus perturbations particularly on the semantically salient regions, effectively fooling the victim classifier. Left: (a) benign input image, (b) generated perturbation (normalized for visual purposes only), (c) unbounded adversarial image, and (d) bounded adversarial image. The label on top (green) and bottom (orange) denote the correct label and prediction after the attack, respectively. Right: We highlight that our method induces Grad-CAM [56] to focus on drastically different regions in our adversarial examples compared to both the benign image and the adversarial examples crafted by the baseline [86]. Moreover, our approach noticeably spreads and reduces the high activation regions observed in the benign and baseline cases, enhancing the transferability of our adversarial perturbations.
Figure S8: Additional cross-task (SS) qualitative results.
@@
Following BIA [86] but expanding to all generator layers, we visualize the difference map in the feature activations between each baseline and ours added to it to better demonstrate the distinct noise our method further adds for enhanced transferability. To compute the difference map at each block, we take the activation map and apply cross-channel average pooling. Then, to better qualitatively highlight the added noise, we further threshold the difference map as follows:
@@
Further looking into how robust our attack performs against purification methods such as NRP [49], we report the improvements with Ours in Table S10. On most of the baselines, our method addition maintains lower Accuracy, ASR and FR scores than the baseline alone, while on the recent advanced methods (e.g. FACL [82] and PDCL [83]), our method very slightly maintains similar scores with those of the baseline alone on the three metrics. On the other hand, on the more challenging ACR metric, we observe that our method slightly improves from the baseline, entailing that our generative feature-level tuning further reduces the number of inadvertently corrected samples.
@@
We also evaluated our method on the zero-shot image classification task with the well-known CLIP [52] vision-language model in Table S10. Here, we observe that, except for BIA [86] and FACL [82], we observe boosted attacked accuracy when we add our method to the baselines [50, 29, 1, 83]. We conjecture that the slight attack strength degradation owes to the baseline method that has already been well-fitted to generate adversarial examples effective for the zero-shot setting based on the relatively lower accuracy scores than the rest. We posit that the respective well-trained generator is already adept enough that our method may interfere with the learned generator weights negatively, and there may exist a maximum capacity at which AEs from generative model-based attacks can attack victim models.
Random trials for Baseline [86] with Ours.
In Table S10, we further show that our method exhibits stable training results (mean±$\pm$±std.dev.) as shown from multiple random seed trials evaluated on all four cross-settings.
@@
Method Metric NRP [49] NRP-ResNet [49]
@@
Table S9: Attack robustness on zero-shot image classification task with CLIP [52].
CLIP ViT-B/16 (%) Acc. ($\Delta$%p) ViT-B/32 (%) Acc. ($\Delta$%p)
Benign 66.10 - 61.02 -
CDA [50] 40.34 - 32.53 -
w/ ours 40.22 -0.12 32.07 -0.46
LTP [29] 40.61 - 33.36 -
w/ ours 40.21 -0.40 33.20 -0.16
BIA [86] 39.27 - 32.16 -
w/ ours 39.78 +0.51 32.70 +0.54
GAMA [1] 40.10 - 33.18 -
w/ ours 39.87 -0.23 32.98 -0.20
FACL [82] 39.50 - 32.69 -
w/ ours 39.63 +0.13 32.92 +0.23
PDCL [83] 40.16 - 33.11 -
w/ ours 39.58 -0.58 32.76 -0.35
Table S10: Random trials for Ours added to the baseline [86].
Trial Cross- Accuracy ($\downarrow$) ASR ($\uparrow$) FR ($\uparrow$) ACR ($\downarrow$)

### v4

@@ C.2 Implementation details
We train the perturbation generator for one epoch on the ImageNet-1K Russakovsky et al. (2015) dataset using the Adam Kingma and Ba (2015) optimizer $\beta=(0.5,0.99)$. We set the learning rate $lr=2e^{-4}$.
@@
Dataset ImageNet-1K Russakovsky et al. (2015) CUB-200-2011 Wah et al. (2011) Stanford Cars Krause et al. (2013) FGVC Aircraft Maji et al. (2013)
@@
Resolution $\times224$ $\times448$ $\times448$ $\times448$
We describe the statistics of the datasets used for training and evaluation in Table C.2. Note that we do not use the training sets from CUB-200-2011 Wah et al. (2011), Stanford Cars Krause et al. (2013), or FGVC Aircraft Maji et al. (2013) for the strict black-box cross-domain setting.
@@
Since our approach only involves computational overheads during the training of the perturbation generator, there is no inference time overhead. During training, we describe in the table on the right that the time for a single forward pass with a batch size of 1 incurs an additional +12.18 (ms) relative to the baseline Zhang et al. (2022) time of 44.72 (ms), and an additional +28.31 (MB) in memory relative to the baseline memory requirement of 1,404.13 (MB), which are averaged over 1,000 iterations (See Table S4). We note that the training was performed on a single NVIDIA RTX A6000 GPU.
@@ C.3 Additional quantitative results
While we performed experiments against the VGG-16 surrogate model for fair comparison with previous works, we also provide, in Tables S5 and S6, our improvements when trained against other surrogate models such as VGG-19, ResNet-152, and DenseNet-169 as practiced in Zhang et al. (2022); Yang et al. (2024a, b). Ours added to the baseline trained against all three surrogate models across models and domains, except for cross-model against Dense169, consistently enhances the attack transferability. As with our results against VGG-16 in Tables 1 and 2, our method effectively boosts the attack capacity regardless of the type of surrogate model used for training the generator. We believe the slight increase in cross-model average when using DenseNet-169 as the surrogate model is driven mainly by pronounced gains on a few architectures (e.g. MNASNet and DeiT) while most other models also benefited, albeit to a lesser extent. In the cross-domain evaluation, adversarial examples generated with DenseNet-169 consistently deliver a substantial average performance boost, underscoring its effectiveness across differing data distributions.
In qualitative comparisons across different surrogate models in Fig. S4, our Grad-CAM Selvaraju et al. (2017) visualizations reveal activation patterns that depart significantly from the baseline. Rather than highlighting only the main object region, our method further amplifies those top responses and draws out additional high-sensitivity areas that the baseline misses. When we step through the stages of adversarial noise generation, we see that our approach consistently places perturbations along object edges and contours. By focusing noise on these shared, model-agnostic features instead of scattering it elsewhere, our method not only seeks to align the generated adversarial noise with the most semantically meaningful regions but also achieves stronger transferability.
Table S5: Additional quantitative cross-model transferability results. We report the average improvement margins of our method added to the baseline Zhang et al. (2022) using different surrogate models (VGG-19, Res152, Dense169). We report the improvements ($\Delta$ %p) with ours relative to the baseline. Better averaged results are marked in boldface.
@@
BIA Zhang et al. (2022) w/ Ours Acc. ($\Delta$ %p) $\downarrow$ -1.49 +0.88 -1.70 -3.49 -4.04 -1.07 +1.06 -2.31 +0.46 -1.23 -4.40 +0.39 -0.24 -0.16 -6.38 -0.17 -0.32 +0.10 -0.55 -2.22 -2.29 -1.72 -1.40
ASR ($\Delta$ %p) $\uparrow$ +1.93 -1.12 +2.17 +4.43 +5.18 +1.34 -1.43 +3.98 -0.65 +1.77 +5.18 -0.44 +0.35 +0.20 +7.63 +0.25 +0.32 -0.24 +0.75 +2.58 +2.75 +1.91 +1.77
FR ($\Delta$ %p) $\uparrow$ +1.72 -0.89 +1.97 +3.95 +4.68 +1.24 -1.19 +2.89 -0.55 +1.42 +4.77 -0.37 +0.42 +0.15 +7.08 +0.09 +0.33 -0.33 +0.72 +2.48 +2.67 +1.95 +1.60
ACR ($\Delta$ %p) $\downarrow$ -0.23 +0.11 -0.35 -0.56 -0.41 -0.13 +0.34 -0.18 -0.20 -0.29 -0.82 0.24 +0.14 -0.02 -1.45 +0.18 -0.33 -0.24 +0.06 -0.93 -0.68 -0.75 -0.30
@@
BIA Zhang et al. (2022) w/ Ours Acc. ($\Delta$ %p) $\downarrow$ -0.61 -0.99 -2.96 -0.38 -5.13 -7.36 -5.35 -0.71 -3.87 -0.89 -6.17 -3.75 -0.12 -0.84 +0.67 -0.94 +0.20 +0.33 -2.37 -4.19 -9.23 -0.46 -2.51
ASR ($\Delta$ %p) $\uparrow$ +0.78 +1.28 +3.76 +0.39 +6.36 +9.09 +7.37 +1.32 +4.75 +1.24 7.16 +4.70 +0.18 +1.13 -0.97 +0.92 -0.24 -0.32 +3.03 +5.09 +11.10 +0.48 +3.12
FR ($\Delta$ %p) $\uparrow$ +0.78 +1.09 +3.23 +0.27 +5.70 +8.14 +6.29 +0.97 +4.37 +0.86 +6.79 +4.18 +0.17 +1.35 -1.04 +1.08 -0.28 -0.55 +2.96 +5.01 +10.83 +0.38 +2.84
ACR ($\Delta$ %p) $\downarrow$ -0.10 +0.01 -0.69 -0.35 -1.21 -1.26 -1.33 +0.06 -0.50 -0.29 -1.56 -0.68 +0.07 +0.19 -0.50 -1.06 +0.02 +0.35 -0.37 -0.96 -2.64 -0.34 -0.60
@@
BIA Zhang et al. (2022) w/ Ours Acc. ($\Delta$ %p) $\downarrow$ -2.25 -7.65 -2.24 -2.76 -12.99 +4.65 +10.36 +2.34 +31.22 -6.50 +1.19 -6.71 -0.62 -0.73 +11.05 +8.35 +0.38 +0.58 +3.63 -2.67 +11.51 -7.61 +1.48
ASR ($\Delta$ %p) $\uparrow$ +2.77 +9.41 +2.77 +3.38 +16.23 -5.34 -16.99 -6.14 -37.55 +9.68 -1.55 +8.22 +0.64 +0.70 -13.29 -9.27 -0.70 -0.77 -4.34 +3.20 -14.21 +8.67 -2.02
FR ($\Delta$ %p) $\uparrow$ +2.41 +8.48 +2.32 +2.88 +13.97 -4.84 -12.84 -4.23 -33.64 +7.93 -1.23 +7.24 +0.76 +0.75 -12.09 -9.48 -0.70 -0.65 -4.29 +2.99 -12.99 +8.44 -1.76
ACR ($\Delta$ %p) $\downarrow$ -0.73 -1.65 -0.71 -0.83 -2.64 +1.17 +1.78 +0.18 +6.26 -0.93 -0.40 -1.75 -0.54 -0.80 +2.23 +4.10 -1.31 +0.16 +1.46 -0.80 +2.00 -2.56 +0.17
Table S6: Additional quantitative cross-domain transferability results. We report the average improvement margins of our method added to each baseline, averaged over three models for each domain using different surrogate models (VGG-19, Res152, Dense169). We report the improvements ($\Delta$ %p) with ours relative to the baseline Zhang et al. (2022). Better averaged results are marked in boldface.
Cross-domain CUB-200-2011 Wah et al. (2011) Stanford Cars Krause et al. (2013) FGVC Aircraft Maji et al. (2013) Avg. Acc. (%)
@@
Figure S4: Comparison of the crafted AEs, Grad-CAMs, intermediate feature activation maps from the baseline, and with ours (columns 1–4, respectively). We present the qualitative results from training against other surrogate models: VGG-19 (Left), ResNet-152 (Center), and DenseNet-169 (Right) as commonly compared in existing generative attacks Zhang et al. (2022); Aich et al. (2022); Yang et al. (2024a, b). The correct label and attacked prediction results are marked in green and white, respectively.
@@
DeepLabV3+ Chen et al. (2018) SegFormer Xie et al. (2021) Faster R-CNN Girshick (2015) DETR Carion et al. (2020)
@@
To further verify that our proposed method is compatible with the baseline Zhang et al. (2022) shown to perform best at the selected mid-level layer of the VGG-16 surrogate model (i.e., Maxpooling.3), we conducted an ablation study on the mid-level layer in Fig. S5. Across domain, model, and two tasks (SS and OD), we observe that our method added to the baseline still maintains the best strength of the attack at the selected mid-level layer, Maxpooling.3, compared to the other early or late layers.
@@ C.4 Additional qualitative results
Figure S7: Additional qualitative results. Our semantic structure-aware attack successfully guides the generator to focus perturbations particularly on the semantically salient regions, effectively fooling the victim classifier. Left: (a) benign input image, (b) generated perturbation (normalized for visual purposes only), (c) unbounded adversarial image, and (d) bounded adversarial image. The label on top (green) and bottom (orange) denote the correct label and prediction after the attack, respectively. Right: We highlight that our method induces Grad-CAM Selvaraju et al. (2017) to focus on drastically different regions in our adversarial examples compared to both the benign image and the adversarial examples crafted by the baseline Zhang et al. (2022). Moreover, our approach noticeably spreads and reduces the high activation regions observed in the benign and baseline cases, enhancing the transferability of our adversarial perturbations.
Figure S8: Additional cross-task ($S$) qualitative results.
@@
Following BIA Zhang et al. (2022) but expanding to all generator layers, we visualize the difference map in the feature activations between each baseline and ours added to it to better demonstrate the distinct noise our method further adds for enhanced transferability. To compute the difference map at each block, we take the activation map and apply cross-channel average pooling. Then, to better qualitatively highlight the added noise, we further threshold the difference map as follows:
@@
Further looking into how robust our attack performs against purification methods such as NRP Naseer et al. (2020), we report the improvements with Ours in Table S11. On most of the baselines, our method addition maintains lower Accuracy, ASR and FR scores than the baseline alone, while on the recent advanced methods (e.g. FACL Yang et al. (2024a) and PDCL Yang et al. (2024b)), our method very slightly maintains similar scores with those of the baseline alone on the three metrics. On the other hand, on the more challenging ACR metric, we observe that our method slightly improves from the baseline, entailing that our generative feature-level tuning further reduces the number of inadvertently corrected samples.
@@
We also evaluated our method on the zero-shot image classification task with the well-known CLIP Radford et al. (2021) vision-language model in Table LABEL:tab:supp_zsclassification. Here, we observe that, except for BIA Zhang et al. (2022) and FACL Yang et al. (2024a), we observe boosted attacked accuracy when we add our method to the baselines Naseer et al. (2019); kanth Nakka and Salzmann (2021); Aich et al. (2022); Yang et al. (2024b). We conjecture that the slight attack strength degradation owes to the baseline method that has already been well-fitted to generate adversarial examples effective for the zero-shot setting based on the relatively lower accuracy scores than the rest. We posit that the respective well-trained generator is already adept enough that our method may interfere with the learned generator weights negatively, and there may exist a maximum capacity at which AEs from generative model-based attacks can attack victim models.
Random trials for Baseline Zhang et al. (2022) with Ours.
[…]
@@
Method Metric NRP Naseer et al. (2020) NRP-ResNet Naseer et al. (2020)
@@
Table S9: Real-world system evaluations (400 random images).
Task Zero-Shot Cls. (Acc. %)
Data subset ImageNet-R ImageNet-1K
Model GPT-4o-mini GPT-4o-mini GPT-4.1
Benign 27.25 34.25 17.50
BIA Zhang et al. (2022) 6.75 5.00 3.00
w/ Ours 4.75 1.25 1.00
Table S10: Multi-modal large language model evaluations with a prompt ‘Provide a short caption for this image‘.
Task Image Captioning (MS COCO) on LLaVA 1.5-7B
Metric BLEU-4 METEOR ROUGE-L CIDEr SPICE
Benign 34.9 29.1 58.0 122.7 22.9
BIA Zhang et al. (2022) 35.0 28.6 57.4 124.9 22.7
w/ Ours 34.5 28.4 57.0 123.1 22.2
Task Image Captioning (MS COCO) on LLaVA 1.5-13B
Metric BLEU-4 METEOR ROUGE-L CIDEr SPICE
Benign 36.9 29.0 58.4 130.6 22.9
BIA Zhang et al. (2022) 36.1 28.6 57.6 127.7 22.6
w/ Ours 36.2 28.6 58.0 128.6 22.6
Table S11: Random trials for Ours added to the baseline Zhang et al. (2022).
Trial Cross- Accuracy $\downarrow$ ASR $\uparrow$ FR $\uparrow$ ACR $\downarrow$